# Faster Convergence on Heterogeneous Federated Edge Learning: An Adaptive Clustered Data Sharing Approach

### v2

@@ IV-B Distribution-based Adaptive Clustering Algorithm
proof. See Appendix Appendix A.
@@ IV-C Joint Computed Frequency and Shared Data Volume Optimization Algorithm
where ${f_{k,\gamma}}$ satisfy the equality $\gamma_{k}^{u}+\varsigma{L_{k}}E{\tilde{n}_{k}}f_{k,\gamma}^{2}={\gamma_{th}}$. When optimizing the shared data volume $\boldsymbol{{N}^{s}}$, the objective of subproblem $\mathcal{P}2$ can be reformulated as follows:
@@
To address the no closed-form challenge for $T\left(\cdot\right)$, we employ data fitting to estimate the uncertain function parameters. Due to the complexity involved in directly fitting the function that incorporates multiple variables ${\boldsymbol{N}^{s}}$, we instead utilize an intermediate function ${\tilde{D}_{{\text{EMD}}}}\left({\boldsymbol{N}^{s}}\right)$ to capture the relationship between $T$ and data quality. Specifically, the number of communication rounds with respect to the data heterogeneity can be expressed as [^1] The detailed proof process can be refered to Appendix B, Theorem 2, and similar results also appear in [11].
@@
5      for $i=0,...,I$ do // Stochastic SCA Algorithm
@@ IV-D Complexity Analysis
Without loss of generality, we discuss the computational complexity of the proposed DACA and JFVO algorithms. When addressing the clustering problem, the predominant computational consumption arises from the generation of the constraint graph. The complexity of calculating these constraint edges in DACA is ${\rm O}\left({{K^{2}}}\right)$. For the joint optimization problem, the computational complexity mainly relies on the calculation of the computed frequency, the construction of a recursive convex approximation function, and the solution of the shared data volume. In the outer iteration, the complexity of updating frequency $\boldsymbol{F}$ is ${\rm O}\left({{JK}}\right)$. During the inner iteration, we design a recursive convex approximation for optimizing the shared data volume. The complexity in calculating the surrogate function is ${\rm O}\left({{JIM}}\right)$, where $M$ is the number of the clusters determined by DACA. Additionally, problem (35) is a quadratic problem (QP) [39] and the complexity of solving it using a standard convex optimization method is denoted by ${\rm O}\left({{JIM^{2}}}\right)$.
In summary, besides the advantages of the DACA discussed in Section IV-B, our proposed algorithm can effectively accomplish clustered data sharing and accelerate FEEL training in uncertain environments. Stochastic optimization algorithms can mitigate the instability resulting from parameter estimation. Moreover, the computational requirement of JFVO is affordable, making it suitable for practical use as an online algorithm.
@@ V Simulation and Experimental Results
Parameter Value Parameter Value
$I$ 3 dB $\gamma_{\text{th}}$ 0.005 J
$R$ 10 $N_{0}$174 dBm/Hz
$P_{\text{BS}}$ 1 W $P_{k}$ 0.01 W
$B^{d}$ 20 MHz $f_{\max}$ $1.2\times 10^{3}$ MHz
$B^{u}$ 1 MHz $L_{k}$ $2.5\times 10^{5}$ cycles/sample
$B_{m}^{s}$ 1 GHz $\varsigma$ $4\times 10^{-26}$
@@ V-A Experimental Setup
Datasets: We evaluate our results using the MNIST [40] and CIFAR-10 datasets [41], which are commonly used image datasets. The MNIST dataset consists of 70,000 grayscale images of handwritten digits from 0 to 9, with each image being a 28x28 pixel square. The dataset is divided into two main sets: a training dataset with 60,000 images and a test dataset with 10,000 images. It is worth noting that each MNIST image consists of 784 pixels, corresponding to $8\times 784$ bits, and a label consisting of 8 bits. Therefore, the total size of each sample is $a=6280$ bits. The CIFAR-10 dataset consists of 10 classes of colored objects with 50,000 training and 10,000 test samples. The local dataset of users are the subsets of the training dataset, and the processing details for these non-IID datasets are as follows:
@@
The feature distribution skew is set by the noise-based feature imbalance [12]. The entire dataset is randomly and equally divided among multiple parties. Then all parts add different levels of Gaussian noise to their local datasets to achieve distinct feature distributions.
2) Model Structure: We adopt Convolutional Neural Networks (CNNs) as they are widely used in image classification tasks. For the MNIST dataset, we employ a $2$-layer CNN with two $5\times 5$ convolutional layers and two fully connected layers. During local training, we set the batch size to $10$, the number of epochs $E$ to $1$, and the learning rate $\eta$ to $0.01$ in the SGD optimizer. For the more complex CIFAR-10 dataset, we employ the famous CNN architecture known as VGG-16 [42]. We set the local epoch $E$ to $5$, the local batch size to $50$, and the learning rate $\eta$ to $0.25$. We compare the proposed method to other algorithms on terms of both the accuracy and the communication cost.
(a) Loss with MNIST data
@@
(c) Loss on Feature Distribution Skew
(d) Accuracy on MNIST data
(e) Accuracy on CIFAR-10 data
(f) Accuracy on Feature Distribution Skew
Figure 4: The performance compared among baselines and ours, ${v_{th}}=3.5\times 10^{5}$ bit/s and ${e_{th}}=0.5$.
@@
Figure 5: The illustration of the number of samples per class allocated to each user (indicated by dot sizes).
@@ V-B Experiment Results
In Fig. 4, we compare the proposed DACA with several baselines. Although DACA is primarily designed for label distribution skew, its applicability can be readily extended to other non-IID cases such as feature distribution skew. The left and middle columns of Fig. 4 show the superior performance of the algorithm on different datasets, and the right column shows DACA’s generalization in the performance of feature distribution skew. To ensure fairness, the same shared data volume is used for different clustering strategies. We record the loss values of the local models on their training dataset to show the algorithm convergence, and validate the accuracy of the global model on the test dataset.
[…]
It can be observed from Fig. 5a that if the data distribution follows IID distribution, the label distribution across users would be uniform. However, as shown in Fig. 5b, certain users may lack specific labels in non-IID case, leading to distribution skewness. In Fig. 5c, the label distribution after multicasting closely approximates an IID distribution. This is because that our proposed clustered data sharing method effectively addresses skewness by filling missing labels for high EMD users.
@@
To obtain the communication rounds function $T\left(\cdot\right)$, we conduct a series of FL experiments involving heterogeneous datasets with different average-EMD in Fig. 6. The number of communication rounds is recorded when the test accuracy of the global model first meets the specified threshold. The acquired data from these experiments are used to estimate the fitted parameters $\boldsymbol{\beta}$ in (31). We also record the final accuracy of the global models which are trained on various average-EMD datasets with a sufficient number of rounds. In order to demonstrate the effect of data fitting, the Normal Mean Squared Error (NMSE) serves as an important indicator, which measures the disparities between estimated and actual values. The estimated rounds with smaller NMSE indicate a high-precision fitted parameters $\boldsymbol{\beta}$, which helps complete the proposed optimization algorithm in Figs. 7 and 8.
@@
Fig. 6 effectively demonstrates the performance of the function curve $T\left(\cdot\right)$ under different accuracy thresholds. This figure depicts that the number of communication rounds proportionally increases with the rise in the average EMD of the dataset. The rate of increase remains relatively gradual initially, but escalates as the dataset’s non-IIDness becomes more pronounced. Additionally, as the desired accuracy threshold becomes more demanding, the number of communication rounds correspondingly rises. These observations are consistent with the preliminary experiments in the Section II-C, and highlight a trade-off between maximizing global model accuracy and minimizing communiction rounds. Furthermore, as depicted in this figure, a discernible relationship exists between accuracy and statistical heterogeneity, and this correlation exhibits an inverse trend in comparison to the communication rounds. Consequently, it is crucial to identify an appropriate average-EMD value for balancing the communication cost and model accuracy, i.e., the optimal shared data volume ${{\boldsymbol{N}}^{s}}$.
Figure 7: Convergence of the JFVO algorithm.
3) Performance Evaluation of the Optimization Algorithm:
In Fig. 7, the convergence performance of the JFVO algorithm is evaluated at various Signal-Noise-Ratio (SNR). SNR is an important parameter that affects the performance and quality of communication systems. We simulate various communication states by varying the SNR to 10dB, 15dB, and 20dB. This allows us to evaluate the effectiveness of JFVO algorithm in these varied communication states. The depicted results illustrate a consistent decrease in the value of $\Psi$ as the iteration index increases, ultimately converging within a span of 6 iterations. This substantiates the rapid convergence capability of the JFVO algorithm. Moreover, it’s observed that as the SNR increases, the value of $\Psi$ also decreases. This trend is attributed to the reduced delay of clustered data sharing FEEL when users have better channel states in the smaller communication coverage area. A channel with a lower SNR has a higher transmission, which makes the total delay of data sharing and model transmission smaller.
Figure 8: Total delay of federated learning with data sharing.
Fig. 8 illustrates the total delay variation with the number of users. Traditional FL defines FL training delay as total delay, which includes model transmission and local processing. For FedAvg with central sharing, total delay includes data collection, sharing, and FL training. In contrast, clustered data sharing FEEL just considers training and sharing time. In Fig. 8, the total delay of the conventional FedAvg algorithm increases as the number of users grows. This is due to the fact that limited communication resources lead to a reduction in transmission rates and an increase in FL training time. Compared with FedAvg, the total delay of clustered data sharing FEEL can be significantly reduced by employing the DACA and JFOV algorithms. This improvement can be attributed to the optimization of computational resources and a reduction in the number of convergence rounds. Notably, the sharing time gradually increases with the growing number of users. This is because of increased interference among cluster heads, which causes a decrease in the multicast rate of data sharing. Furthermore, the latency of clustered data sharing is slightly lower than that of central sharing, which involves extra delay for data uploading, referred to as Collecting Time.
VI Conclusion
@@
In this work, we investigate and quantify the statistical heterogeneity which is identified as a primary impediment to the convergence of FL. Furthermore, we introduce a clustered data sharing framework to address the non-IID challenge in FEEL and formulate an optimization problem which is aimed at minimize the overall delay in transmission and computation. We decompose the original problem into two subproblem, each dedicated to identifying the optimal clustering strategy and efficient resource allocation. Three conditions are established to aid in clustering process, rooted in a privacy-preserving constrained graph. To enhance the efficacy of data sharing, a clustering algorithm is proposed for selecting cluster heads and credible associates based on the data distribution and the constraints. We derive a upper bound on FEEL convergence to fit the relationship between data distribution and the communication rounds. Then a stochastic optimization approach is employed to determine the optimal computed frequency and shared data volume. The experimental results show that the proposed framework is efficacious in mitigating the data heterogeneity within the system and expediting the FEEL training process. Moreover, our method exhibits well performance not only in resource-constrained environments but also in different non-IID cases.

### v3

@@ IV-B Distribution-based Adaptive Clustering Algorithm
The proof is omitted in this paper due to the limited space. The full proofs can refer to Appendix or online version [40].
@@ IV-C Joint Computed Frequency and Shared Data Volume Optimization Algorithm
where ${f_{k,\gamma}}$ satisfies the equality $\gamma_{k}^{u}+\varsigma{L_{k}}E{\tilde{n}_{k}}f_{k,\gamma}^{2}={\gamma_{th}}$. When optimizing the shared data volume $\boldsymbol{{N}^{s}}$, the objective of subproblem $\mathcal{P}2$ can be reformulated as follows:
@@
To address the no closed-form challenge for $T\left(\cdot\right)$, we employ data fitting to estimate the uncertain function parameters. Due to the complexity involved in directly fitting the function that incorporates multiple variables ${\boldsymbol{N}^{s}}$, we instead utilize an intermediate function ${\tilde{D}_{{\text{EMD}}}}\left({\boldsymbol{N}^{s}}\right)$ to capture the relationship between $T$ and data quality. Specifically, the number of communication rounds with respect to the data heterogeneity can be expressed as [^3] The detailed proof process can be refered to Appendix B, Theorem 2 or online version [40], and similar results also appear in [11].
@@
5      for $i=0,...,I$ do // Inner iteration Stochastic SCA Algorithm
@@ IV-D Complexity Analysis
Without loss of generality, we discuss the computational complexity of the proposed DACA and JFVO algorithms. When addressing the clustering problem, the predominant computational consumption arises from the generation of the constraint graph. The complexity of calculating these constraint edges in DACA is ${\rm O}\left({{K^{2}}}\right)$. For the joint optimization problem, the computational complexity mainly relies on the calculation of the computed frequency, the construction of a recursive convex approximation function, and the solution of the shared data volume. In the outer iteration, the complexity of updating frequency $\boldsymbol{F}$ is ${\rm O}\left({{JK}}\right)$. During the inner iteration, we design a recursive convex approximation for optimizing the shared data volume. The complexity in calculating the surrogate function is ${\rm O}\left({{JIM}}\right)$, where $M$ is the cluster number determined by DACA. Problem (35) is a quadratic problem (QP) [42] and the complexity of solving it using a standard convex optimization method is denoted by ${\rm O}\left({{JIM^{2}}}\right)$.
In summary, besides the advantages of the DACA discussed in Section IV-B, our proposed algorithm can effectively accomplish users group and accelerate FEL training in uncertain environments. Stochastic optimization algorithms can mitigate the instability resulting from parameter estimation. Moreover, the computational requirement of DACA and JFVO is affordable, making it suitable for practical use as an online algorithm.
@@ V Simulation and Experimental Results
Parameter Value
Multicast interference $I$ 3 dB
Number of subcarriers $R$ 10
Nosie power spectral density $N_{0}$174 dBm/Hz
Transmit power of the BS $P_{\text{BS}}$ 1 W
Transmit power of user $k$ $P_{k}$ 0.01 W
Bandwidth of multicasting $B_{m}^{s}$ 1 GHz
Bandwidth of broadcating $B^{d}$ 20 MHz
Bandwidth of uploading $B^{u}$ 1 MHz
Energy consumption requirement $\gamma_{\text{th}}$ 0.005 J
Energy consumption coefficient $\varsigma$ $4\times 10^{-26}$
Maximum computation frequency $f_{\max}$ $1.2\times 10^{3}$ MHz
CPU cycles for per sample $L_{k}$ $2.5\times 10^{5}$ cycles/sample
@@ V-A Experimental Setup
Datasets: We evaluate our results using the MNIST [43] , CIFAR-10 datasets [44] and Shakespeare [45], which are commonly used image classification and next word prediction. The MNIST dataset consists of a training dataset with 60,000 images and a test dataset with 10,000 images. The CIFAR-10 dataset consists of 10 classes of colored objects with 50,000 training and 10,000 test samples. Shakespeare contains text dataset from the complete works of William Shakespeare and dialogue of each character is used as one dataset on a device. Naturally, Shakespeare dataset exhibits feature distribution skew. The processing details for these non-IID datasets are as follows:
@@
The feature distribution skew: is set by the noise-based feature imbalance [12]. The entire dataset is randomly and equally divided among multiple parties. Then all parts add different levels of Gaussian noise to their local datasets to achieve distinct feature distributions. Shakespeare dataset on devices already have different feature distributions and does not need to be preprocessed.
2) Model Structure: We adopt Convolutional Neural Networks (CNNs) as they are widely used in image classification tasks. For the MNIST dataset, we employ a $2$-layer CNN with two $5\times 5$ convolutional layers and two fully connected layers. For the more complex CIFAR-10 dataset, we employ the famous CNN architecture known as VGG-16 [46]. For the Shakespeare dataset, we adapt a Long Short-Term Memory (LSTM) model as the network architecture of NLP task.
(a) Loss on MNIST data
@@
(c) Loss on Shakespeare data
(d) Loss on Noise-based data
(e) Accuracy on MNIST data
(f) Accuracy on CIFAR-10 data
(g) Accuracy on Shakespeare data
(h) Accuracy on Noise-based data
Figure 4: The convergence and accuracy performance with different data sharing strategies, ${v_{th}}=3.5\times 10^{5}$ bit/s and ${e_{th}}=0.5$.
@@
Figure 5: The illustration of the sample number per class allocated to each user (indicated by dot sizes).
@@
To demonstrate the superiority of the resource allocation algorithm, the JFVO algorithm is compared with some existing SOTA (State-of-the-Art) works:
  • The stochastic Majorization-Minimization (MM) [48]: This algorithm focuses on constructing a surrogate function that majorizes (upper bounds) the original objective function and iteratively minimizes this surrogate function, ensuring convergence by reducing the original objective at each step.
  • Sample Average Approximation (SAA) [27]: This method can solve stochastic optimization problems by replacing the expectation in objective function with sample average. And the stochastic problem is transformed into a deterministic one that can be solved by standard optimization techniques.
@@ V-B Experiment Results
In Fig. 4, we compare the proposed DACA with several baselines. Although DACA is primarily designed for label distribution skew, its applicability can be readily extended to other non-IID cases such as feature distribution skew. The left two columns of Fig. 4 show the superior performance of the algorithm on label distribution skew, and the right columns shows DACA’s generalization in the performance of feature distribution skew. Moreover, we conduct experiments on the commonly used Dirichlet distribution. The results in Table II confirm that our method remains effective under this setting.
From the top row of Fig. 4, it is evident that the proposed clustered data sharing methods outperform others in terms of reducing the number of communication rounds and achieving lower loss values. Additionally, the bottom of Fig. 4 demonstrate that the proposed framework significantly enhances test accuracy. The CIFAR-10 dataset, being more complex in feature distribution compared with MNIST dataset, exhibits more pronounced improvements. This demonstrates that our method can achieve greater performance gains for training with initially poor accuracy. The DACA shows favorable results even compared with central sharing method on CIFAR-10 and noise-based dataset. This is because central sharing only broadcast the same samples to all participants while clustered sharing multicast different samples to augment local dataset. Moreover, the DACA algorithm improves the accuracy by 6% on the Shakespeare dataset, which highlights the applicability of the proposed framework to natural language processing (NLP) tasks. These observation demonstrate the superiority of DACA over other clustering strategy in terms of estimating statistical heterogeneity.
It can be observed from Fig. 5(a) that if the data distribution follows IID distribution, the label distribution across users would be uniform. However, as shown in Fig. 5(b), certain users may lack specific labels in non-IID case, leading to distribution skewness. In Fig. 5(c), the label distribution after multicasting closely approximates an IID distribution. This is because that our proposed clustered data sharing method effectively addresses skewness by filling missing labels for high EMD users.
TABLE II: Comparison of performance on Dirichlet distribution.
Dataset Dirichlet FedAvg SCC CEC Central DACA(ours)
MNIST Dir=0.1 96.17 97.42 97.50 97.84 97.98$(\uparrow 1.8)$
Dir=1.0 97.45 97.86 97.90 98.07 98.09$(\uparrow 0.6)$
Dir=10 97.80 97.90 97.94 97.97 98.08$(\uparrow 0.2)$
CIFAR-10 Dir=0.1 67.72 74.00 74.93 75.73 75.90$(\uparrow 8.2)$
Dir=1.0 75.10 75.88 76.05 76.65 76.48$(\uparrow 1.3)$
Dir=10 76.70 77.93 77.95 78.04 77.54$(\uparrow 0.8)$
@@
To obtain the communication rounds function $T\left(\cdot\right)$, we conduct a series of FL experiments involving heterogeneous datasets with different average-EMD in Fig. 6. The number of communication rounds is recorded when the test accuracy of the global model first meets the specified threshold. The acquired data from these experiments are used to estimate the fitted parameters $\boldsymbol{\beta}$ in (31). We also record the final accuracy of the global models which are trained on various average-EMD datasets with a sufficient number of rounds. In order to demonstrate the effect of data fitting, the Normal Mean Squared Error (NMSE) serves as an important indicator, which measures the disparities between estimated and actual values. The predicted values have low NMSE which indicates the high-precision of fitted round-EMD function.
@@
Fig. 6 effectively demonstrates the performance of the function $T\left(\cdot\right)$ under different accuracy thresholds. This figure depicts that as the average EMD rises, the number of communication rounds proportionally increases. The rate of increase relatively gradual initially, but then escalates exponentially. Additionally, as the desired accuracy threshold becomes more demanding, the number of communication rounds correspondingly rises. These observations are consistent with the preliminary experiments in the Section II-C, and highlight a trade-off between maximizing global model accuracy and minimizing communication rounds. Furthermore, As shown in the blue line, a discernible relationship exists between accuracy and statistical heterogeneity, and this correlation exhibits an inverse trend in comparison to the communication rounds. Consequently, it is crucial to identify an appropriate average-EMD value for balancing the communication cost and model accuracy, i.e., the optimal shared data volume ${{\boldsymbol{N}}^{s}}$.
(a) Convergence versus resource allocation algorithms.
(b) Convergence versus SNR.
Figure 7: Convergence of the JFVO algorithm. Average Sharing Ratio (ASR) is the weighted average of local sharing ratio of cluster heads, which measures the degree of data leakage. $ASR=\sum\limits_{m\in\mathcal{M}}{\frac{{\left|{{\mathcal{C}_{m}}}\right|}}{K}\cdot\frac{{n_{m}^{s}}}{{{n_{m}}}}}$.
3) Convergence of the JFVO Algorithm:
In Fig. 7, we compare the convergence of the proposed JFVO algorithms with traditional approaches and verify its robustness in different Signal-Noise-Ratio (SNR) environments. As shown in Fig. 7(a), JFVO consistently outperforms baseline methods in both convergence speed and optimization results, showcasing its ability to effectively manage trade-offs between resource allocation and FL training. This superior performance can be attributed to JFVO’s effectiveness in tackling stochastic non-convex objectives, making it a practical and scalable solution for heterogeneous wireless FL environments.
Furthermore, we evaluate the performance of the JFVO algorithm under SNRs of 10 dB, 15 dB, and 20 dB. As shown in Fig. 7(b), the total training delay keeps decreasing as the iteration index increases, and ultimately converges within 6 iterations. Therefore, it substantiates the rapid convergence capability of the JFVO algorithm. Additionally, we also record the shared data volume as the algorithm converges, and the results show that the proposed framework can achieve significant performance gains of delay with little data leakage (less than 1.0). Interestingly, as the SNR of downloading increases, the shared data volume also increases leading to more privacy leakage. This is because that download delay at low SNR is high which requires more shared data volume to reduce the number of communication rounds.
Figure 8: The complex comparison of resource allocation algorithms.
4) Complexity Comparison of Optimization Algorithms:
In addition to the computational complexity analysis, we evaluate the algorithm complexity is evaluated based on CPU running time, which are measured on the same machine (Intel Xeon Gold 6139M). For fairness, the start and end times of each process are recorded to compare the computational efficiency of JFVO against traditional approaches. As illustrated in Fig. 8, JFVO demonstrates low and stable computational overhead compared to baseline algorithms. This efficiency is attributed to its recursive convex approximation function, which enables rapid convergence to optimal solutions while significantly reducing inter-iteration times. Furthermore, as the number of users increases, the running times of the SAA and SMM algorithms grow significantly, while the running time of JFVO remains nearly constant. This result shows superior capability of JFVO in handling large-scale optimization problems effectively.
(a) Shared data volume and SINR of each cluster
(b) Total delay of FL versus number of users.
Figure 9: Resource allocation of clustered data sharing FL
5) Resource Allocation of the Proposed Framework:
We record the shared data volume for each cluster head and the SINR within the clusters. As shown in Fig. 9(a), the sharing data ratio of the cluster head decreases as the multicasting SINR diminishes. This is because the delay of data sharing is determined by the cluster with the lowest multicast rate, allowing clusters with higher SINR to exchange more data within the same multicast delay. Compared with other algorithms, JFVO can allocate more shared data in each cluster, which reduces statistical heterogeneity more effectively in FL.
The Fig. 9(b) provides a detailed analysis for delay of proposed framework on communication and computation. As shown in Fig. 9(b), the proposed JFVO algorithm outperforms both SMM and SAA in minimizing the total delay under resource-constrained conditions. This is due to its ability to effectively balance the trade-off between muticasting and FEL training latencies by jointly optimizing shared data volume and computed frequency. Notably, the shared data time of JFVO slightly increases as the number of users grows. This result is caused by increased interference among cluster heads, which leads to a lower multicast rate of data sharing. The FEL training delay (including broadcasting, computation and uploading) is less than other algorithms. This improvement stems from the reduction in the number of convergence rounds required for FEL.
VI Conclusion and Future Work
@@
In this work, we introduce a clustered data sharing framework to address the non-IID challenge in FEL and formulate an optimization problem which is aimed at minimize the overall delay in transmission and computation. We decompose the original problem into two subproblem, each dedicated to identifying the optimal clustering strategy and efficient resource allocation. Three conditions are established to aid in clustering process, rooted in a privacy-preserving constrained graph. To enhance the efficacy of data sharing, a clustering algorithm is proposed for selecting cluster heads and credible associates based on the data distribution and the constraints. We derive a upper bound on FEL convergence to fit the relationship between data distribution and the communication rounds. Then a stochastic optimization approach is employed to determine the optimal computed frequency and shared data volume. The experimental results show that our method exhibits well performance not only in resource-constrained environments but also in different non-IID cases.
To further enhance model accuracy, we propose the following two potential directions: i) Efficiently share high-quality synthetic data: Sharing more high-quality data can improve accuracy further, However, this approach may trade off communication efficiency for enhanced model performance. ii) Integrate model-based approaches: Combining our framework with techniques like FedProx or SCAFFOLD can further address heterogeneity, complementing our approach.
Acknowledgment
This work was supported in part by the National Key R&D Program of China (No. 2021YFB3300100), and the National Natural Science Foundation of China (No. 62171062) and (U24A20234).